# Distributional Balancing for Causal Inference: A Unified Framework via Characteristic Function Distance

### v1

@@ 6 Simulation
Distributional Balancing Example 1 Gaussian 100 3.0 63.5 97.8 94.4 4.6 2.7 -3.5 69.9 96.6 93.0 5.1 2.9
200 -1.6 43.1 98.2 95.2 2.4 1.6 -0.6 44.5 97.8 93.4 2.6 1.7
400 1.4 30.8 98.6 96.2 1.5 1.1 1.0 30.7 97.8 96.0 1.6 1.1
800 -1.6 20.4 98.8 98.2 1.0 0.8 -1.0 23.0 97.2 95.6 1.0 0.8
1600 -0.3 16.0 97.2 97.2 0.6 0.6 -0.9 15.2 98.6 98.4 0.7 0.6
Example 2 $\ell_{1}$-Laplacian 100 2.2 68.7 98.0 91.0 5.0 2.5 -12.4 72.9 95.4 88.4 5.6 2.8
200 -1.4 43.5 96.8 91.6 2.6 1.6 -8.6 46.6 96.2 89.0 2.8 1.6
400 0.5 27.8 98.4 92.4 1.6 1.0 -2.1 29.8 96.2 91.4 1.6 1.1
800 -0.8 19.3 97.8 90.6 1.0 0.7 -1.3 20.4 96.8 89.8 1.0 0.7
1600 -0.3 13.3 97.6 91.2 0.7 0.4 -1.6 12.7 99.0 92.0 0.7 0.5
Example 2 $t_{5}$ 100 1.1 71.9 99.6 96.8 6.2 3.5 -2.5 77.2 98.8 96.0 6.5 3.9
200 -2.1 49.2 99.2 96.4 3.2 2.0 2.5 51.4 98.6 95.6 3.3 2.0
400 0.8 34.1 99.6 97.6 1.8 1.3 3.2 34.1 99.4 97.4 1.9 1.3
800 -1.3 20.8 99.6 97.6 1.1 0.9 1.8 22.7 99.4 96.8 1.2 0.9
1600 -0.1 13.8 98.6 96.4 0.7 0.6 0.3 13.8 99.6 97.8 0.7 0.6
Example 3 Matérn with $s=7.5$ 100 0.5 64.3 97.5 91.7 4.6 2.6 -7.1 66.3 95.5 92.1 5.0 2.8
200 -0.0 41.1 97.5 93.9 2.4 1.5 0.3 42.5 97.1 93.3 2.5 1.6
400 1.1 28.4 98.4 94.4 1.5 1.0 1.1 28.6 98.7 93.3 1.5 1.0
800 -1.1 18.6 98.4 96.2 0.9 0.7 -0.4 20.1 98.4 94.2 1.0 0.7
1600 -0.9 13.5 97.8 94.4 0.6 0.5 -0.6 13.4 98.9 95.3 0.6 0.5
Example 4 Energy 100 4.0 62.6 98.0 93.0 4.4 2.4 -10.6 68.4 95.0 89.2 4.8 2.6
200 -0.8 40.0 97.8 91.2 2.4 1.5 -7.1 43.0 95.6 89.4 2.5 1.5
400 0.5 25.4 99.0 91.2 1.5 0.9 -2.5 27.0 97.8 90.6 1.5 0.9
800 -1.1 17.3 98.0 95.2 0.9 0.6 -1.9 18.7 97.6 88.4 1.0 0.6
1600 -0.4 11.6 98.2 90.4 0.6 0.4 -1.5 11.8 98.8 90.4 0.6 0.4
IPW 100 9.4 109.3 - 99.0 - 7.8 -36.8 81.9 - 95.0 - 5.8
200 -1.0 53.4 - 95.6 - 2.4 -32.5 51.4 - 91.8 - 2.2
400 -0.1 34.6 - 95.8 - 1.5 -32.1 38.0 - 83.2 - 1.4
800 0.5 26.4 - 92.8 - 1.0 -32.6 25.5 - 72.2 - 1.0
1600 -0.0 17.4 - 94.8 - 0.7 -33.0 17.3 - 52.4 - 0.7
CBPS 100 4.9 75.2 - 96.2 - 3.7 -33.2 81.3 - 93.2 - 3.9
200 -1.2 51.6 - 95.0 - 2.1 -30.2 51.8 - 91.6 - 2.2
400 -0.1 33.7 - 96.0 - 1.4 -30.9 38.1 - 84.0 - 1.4
800 0.6 26.1 - 92.6 - 1.0 -32.1 25.5 - 74.2 - 1.0
1600 -0.1 17.3 - 94.8 - 0.7 -32.7 17.3 - 54.0 - 0.7
@@ 7 Application
Distributional Balancing Example 1 Gaussian 11927 SS: (7153, 16711) 9558
Boot: (7561, 14679) 7118
Example 2 $\ell_{1}$-Laplacian 12486 SS: (8996, 17261) 8265
Boot: (8470, 14763) 6293
Example 2 $t_{5}$ 12828 SS: (8970, 16979) 8009
Boot: (8345, 15073) 6728
Example 3 Matérn with $s=4.5$ 12307 SS: (7511, 17158) 9648
Boot: (7887, 15012) 7126
Example 4 Energy 12619 SS: (9487, 16997) 7511
Boot: (8206, 14133) 5927
IPW 444 (-12426, 9452) 21878
CBPS 5769 (-486, 11461) 11948
@@ 8 Discussion
We conclude the paper by outlining potential directions for future research. First, the proposed CFD-based method requires the investigator to specify the density $\omega$ and its associated hyperparameters. We have not formally addressed how to select these quantities optimally based on the performance of the resulting weighting estimator. Developing a formal, data-driven cross-validation procedure for selecting $\omega$ and its hyperparameters by directly optimizing the causal effect estimator remains an important open direction. Second, the proposed estimator requires solving a constrained QP involving a gram matrix of size $n\times n$. For large $n$, this computation can become burdensome, motivating the development of more efficient algorithms for exceptionally large datasets. We believe that approximation techniques (e.g., Nyström method or sampling-based approaches) are a promising direction for future research. Finally, in Section 6, we find that our subsampling confidence intervals are generally conservative. Future research could explore ways to refine subsampling procedures to achieve nominal coverage without being overly conservative, or to develop inferential methods that are theoretically valid for estimators derived from constrained optimization problems.

### v2

@@ 6 Simulation
Distributional Balancing Example 1 Gaussian 100 0.9 62.6 97.8 94.0 4.5 2.6 -5.7 67.6 95.8 92.0 5.0 2.8
200 0.9 43.4 98.0 95.4 2.4 1.6 1.2 44.3 98.0 94.0 2.5 1.6
400 1.2 30.5 98.0 96.4 1.5 1.1 1.2 30.2 98.2 95.8 1.5 1.1
800 -1.3 20.4 98.6 98.2 1.0 0.8 -0.2 22.7 97.8 96.2 1.0 0.8
1600 -0.4 16.0 96.4 96.6 0.6 0.6 -0.5 15.2 98.6 99.2 0.7 0.6
Example 2 $\ell_{1}$-Laplacian 100 1.9 65.5 97.8 90.8 5.0 2.5 -13.1 72.5 94.8 88.0 5.7 2.8
200 0.4 42.2 97.2 91.8 2.6 1.6 -6.2 46.2 96.8 89.8 2.8 1.6
400 0.9 28.1 98.0 91.4 1.5 1.0 -1.5 30.5 95.8 91.4 1.6 1.0
800 -0.7 19.7 97.8 89.4 1.0 0.7 -1.4 20.0 97.2 90.2 1.0 0.7
1600 -0.4 13.2 98.2 90.4 0.7 0.4 -1.5 12.8 99.0 91.8 0.7 0.5
Example 2 $t_{5}$ 100 0.2 70.9 99.6 96.4 6.2 3.5 -3.6 76.3 98.4 95.4 6.5 3.7
200 1.1 49.3 99.2 96.6 3.2 2.0 4.0 52.1 99.2 96.2 3.3 2.0
400 0.5 33.1 99.6 97.6 1.8 1.3 3.6 33.6 99.4 97.8 1.9 1.3
800 -0.6 20.7 99.4 97.8 1.1 0.9 2.5 22.9 99.4 97.0 1.1 0.9
1600 -0.2 13.9 98.8 96.6 0.7 0.6 0.5 13.7 99.8 98.0 0.7 0.6
Example 3 Matérn with $s=7.5$ 100 1.2 61.5 98.0 93.4 4.5 2.6 -6.5 66.7 95.8 91.2 5.1 2.8
200 0.7 41.2 97.8 94.4 2.4 1.5 0.4 42.5 97.2 93.4 2.5 1.6
400 1.4 28.3 98.4 94.2 1.5 1.0 0.8 28.2 98.6 93.8 1.5 1.0
800 -1.1 18.2 98.8 96.4 0.9 0.7 -0.4 20.1 98.4 94.2 1.0 0.7
1600 -0.8 13.5 97.6 94.2 0.6 0.5 -0.7 13.2 99.4 95.4 0.6 0.5
Example 4 Energy 100 2.1 62.4 97.8 92.8 4.4 2.4 -12.0 68.6 94.6 88.8 4.9 2.6
200 0.9 39.2 97.6 91.4 2.4 1.5 -4.7 42.4 96.0 90.2 2.5 1.5
400 1.0 25.5 98.8 90.4 1.4 0.9 -2.1 27.5 97.6 90.0 1.5 0.9
800 -0.8 17.2 98.0 89.2 0.9 0.6 -1.6 18.3 98.0 88.4 1.0 0.6
1600 -0.7 11.8 98.0 89.2 0.6 0.4 -1.4 11.8 98.8 90.2 0.6 0.4
IPW 100 8.5 103.3 - 99.0 - 7.6 -35.4 80.2 - 95.0 - 5.8
200 0.2 52.4 - 96.0 - 2.5 -33.4 53.0 - 91.0 - 2.2
400 0.0 34.2 - 96.0 - 1.5 -32.3 38.0 - 83.0 - 1.4
800 -0.3 26.5 - 92.8 - 1.0 -33.2 25.7 - 71.6 - 1.0
1600 -0.5 17.5 - 94.6 - 0.7 -33.1 17.0 - 52.8 - 0.7
CBPS 100 5.4 74.4 - 96.6 - 3.7 -31.7 79.8 - 93.2 - 3.8
200 -0.3 50.6 - 95.6 - 2.1 -31.0 53.3 - 90.8 - 2.2
400 -0.1 33.4 - 96.2 - 1.4 -31.2 38.1 - 84.0 - 1.4
800 -0.2 26.2 - 92.6 - 1.0 -32.6 25.7 - 73.6 - 1.0
1600 -0.5 17.4 - 94.6 - 0.7 -32.8 17.0 - 54.4 - 0.7
@@ 7 Application
Distributional Balancing Example 1 Gaussian 11927 SS: (7153, 17140) 9988
Boot: (7405, 14792) 7387
Example 2 $\ell_{1}$-Laplacian 12486 SS: (8984, 16625) 7641
Boot: (8330, 15076) 6746
Example 2 $t_{5}$ 12828 SS: (7813, 17641) 9828
Boot: (8135, 15126) 6990
Example 3 Matérn with $s=4.5$ 12307 SS: (7613, 17234) 9621
Boot: (7887, 15251) 7364
Example 4 Energy 12619 SS: (9399, 16802) 7404
Boot: (8206, 14359) 6153
IPW 444 (-12648, 9684) 22333
CBPS 5769 (-631, 12024) 12655
@@ 8 Discussion
We conclude the paper by outlining potential directions for future research. First, the proposed CFD-based method requires the investigator to specify the density $\omega$ and its associated hyperparameters. We have not addressed how to select these quantities optimally based on the performance of the resulting weighting estimator. Developing a data-driven cross-validation procedure for selecting $\omega$ and its hyperparameters by directly optimizing the causal effect estimator remains an important open direction. Second, the proposed estimator requires solving a constrained QP involving a gram matrix of size $n\times n$. For large $n$, this computation can be burdensome, motivating the development of more efficient algorithms for exceptionally large datasets. We believe that approximation techniques (e.g., Nyström method or sampling-based approaches) are promising directions for future research. Finally, in Section 6, we find that our subsampling confidence intervals are generally conservative. Future research could explore ways to refine subsampling procedures to achieve nominal coverage without being overly conservative, or to develop inferential methods that are theoretically valid for estimators derived from constrained optimization problems.
@@
Data Availability Statement: The data that support the findings of this study are openly available in DoubleML at https://docs.doubleml.org/stable/index.html.